(a) Derive the demand function of each individual for commodity Q, with the values a = 100,
b1 = 1, b2 = 2, and b3 = 2. Which of the consumers experience diminishing marginal utilities
from the consumption of commodity Q?
1
Expert's answer
2020-09-02T11:37:21-0400
"Q_1=100-p"


"Q_2=100-2p"

"Q_3=100-2p"

The second and third consumers experience a decrease in marginal utility.